Two angles are complimentary. the sum of the measures of the first angle and half the second angle is 69.5 find the measures of the angles
what is the measure of the smaller angle
what is the measure of the other angle

Answer by stanbon(75887)

Two angles are complimentary.
x + y = 90 degrees
-----------
The sum of the measures of the first angle and half the second angle is 69.5
x + (1/2)y = 69.5
-----------
find the measures of the angles
90-y + (1/2)y = 69.5
(-1/2)y = -20.5
y= 41 degrees
x + 41 = 90
x = 49 degrees
--------------------
what is the measure of the smaller angle?: 41 degrees
what is the measure of the other angle?: 49 degrees
--------------------
Cheers,
Stan H.